A rubber band has a force constant of 45 N/m. When the rubber band is pulled by a force of 9.8 N, how far will it stretch and what is its potential energy?


1
Expert's answer
2021-01-19T07:08:44-0500

The elongation of the band is given by the Hook's law:


"x = \dfrac{F}{k}"

where "k = 45N\/m" is the force constant, and "F = 9.8N" is the applied force. Thus, obtain:


"x = \dfrac{9.8N}{45N/m} \approx 0.22m"

The potential energy is given as follows:


"W = \dfrac{kx^2}{2}\\
W = \dfrac{45\cdot 0.22^2}{2} \approx 1.1J"

Answer. 0.22 m, and 1.1J.